OpenStudy (anonymous):

A projectile is thrown upward so that its distance above the ground after t seconds is given by the function h(t) = - 16t^2 + 704t. After how many seconds does the projectile take to reach its maximum height?

OpenStudy (lgbasallote):

do you know differential calculus? specifically derivatives?

OpenStudy (anonymous):

Derivative: h'(t)=-32t+704 h'(t)=0 then: t=704/32 t=22

OpenStudy (anonymous):

where did you get the -32?

OpenStudy (anonymous):

Can't you just find the vertex of the parabola?

OpenStudy (lgbasallote):

derivative of -16t^2..power rule

OpenStudy (lgbasallote):

\[\frac{d}{dx} -16t^2 \implies (2 \times -16)t ^{2-1} \implies -32t\]

OpenStudy (lgbasallote):

then \[\frac{d}{dx} 704t \implies (1 \times 704)t^{1-1} \implies 704t^0 \implies 704\]

OpenStudy (lgbasallote):

so you get \[-32t + 704 = 0\] \[704 = 32t\] \[\frac{704}{32} = \frac{32t}{32}\] \[\frac{704}{32} = t\]

OpenStudy (anonymous):

ohhh i think i get it thanks!

OpenStudy (lgbasallote):

you're welcome and glad you got it :)